HOME ESR EXPERIMENTS LABS etc
E121 E127 E132 E125 E143 laser_cooling_2021 E142 Ê128 E146 E0052 E0018 E0028 E0038 E0075 E0203_E0018
  masses & isomeres 2025  Not logged in ELOG logo
Message ID: 77     Entry time: Wed Apr 2 11:09:42 2025
Author: Michael, Alex 
Category: Accelerator 
Subject: The beam is back in the ESR 
We have the beam again the ESR since around 11:00.
The Schottky spectra look the same as before.
The ESR experts could not exactly identify the problem which caused this loss of the beam.
ELOG V3.1.5-fc6679b